OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. show all work 5/7-5/51

OpenStudy (anonymous):

find commun denominator first: 5/7-5/51 = (5*51 - 5*7)/51*7=?

OpenStudy (anonymous):

(((5 * 51) - (5 * 7)) / 51) * 7 = 30.1960784

OpenStudy (anonymous):

5/7-5/51 = (5*51 - 5*7)/51*7=30 10/51

OpenStudy (anonymous):

come on its to late giv a medal to deserving

Parth (parthkohli):

Hi Rohan, welcome to OpenStudy. Try not to ask for medals here, please. :)

OpenStudy (anonymous):

hello parth how're you??

Parth (parthkohli):

I am good but it's not allowing me to message you.

OpenStudy (anonymous):

how u r studies

Parth (parthkohli):

Wait..\( \color{Black}{\Rightarrow \Large {5 \over 7} - {5 \over 51} = {5(51) - 5(7) \over 51 \cdot 7} }\)